Mai Tai laser

About:

The Mai Tai laser (made by Spectra-Physics, a division of Newport) is primarily operated by the Prairie Technologies PrairieView software.

Our Mai Tai's serial number: 4705 4571-R1 (Specifications on Arrival)

Manual link: Newport / Spectra-Physics HP Mai Tai laser manual (2-photon laser)

Operation:

The laser has a chiller with three settings that you can toggle through, distinguished by what the symbol you see in the left of the little panel on the front of the box.

* = Standby mode. No temperature control.

+ = Heating mode. The box is controlling the temperature.

- = Cooling mode. The box is controlling the temperature.

Leave the laser controller and chiller powered ON and COOLING or HEATING at all times, even when no one is in the lab, even over long breaks, even when everyone is on vacation. It keeps the laser components at safe temperatures that will prolong its life. [Okay, in the event of a planned power outage in the building, turn the laser and chiller off ahead of time by switching the "rocker" power switch to the OFF position.]

The rest of its operation is controlled by the PrairieView program (via a serial connection).

Safety:

Go here for basic laser safety. (add more details here?)

To quote the manual:

At any point during operation, the laser emission can be immediately stopped by pressing the OFF button on the GUI Main screen or by turning the keyswitch to the off (vertical) position.

Maintenance:

If power isn't as expected, try power optimization procedure:

    1. Use the SpectraPhysics/Newport software on the Prairie computer to control the laser.

    2. Turn on "emission" mode (THIS "BUTTON" NEEDS TO BE HELD WITH THE MOUSE FOR 5 SECONDS), no need to open the shutter.

    3. Let the laser run for awhile (0.5 hours) if it hasn't already been running.

    4. Under the Diagnostics panel, click Start AutoTune.

    5. Let it finish!

Every 6 months, we should examine the status of the humidity filter.

    1. Use the SpectraPhysics/Newport software on the Prairie computer to control the laser.

    2. Turn on "emission" mode (THIS "BUTTON" NEEDS TO BE HELD WITH THE MOUSE FOR 5 SECONDS), no need to open shutter.

    3. Under the Diagnostics panel look at the internal humidity reading.

    4. If the humidity reading is greater than 10% order a replacement.

    5. Update the log on the wall (Word document)

Every 6 months, one must change the chiller water and check the chiller filter. (The chiller circulates cold water around the laser to keep it from overheating.) To do this:

    1. Make sure laser is not emitting; that is, it is in stand-by mode (still powered on).

    2. Wheel out the chiller and controller. Prepare a bucket or other reservoir.

    3. Remove the red (labeled "DRAIN HERE") or black quick disconnect tubing, allowing water to flow into the reservoir.

    4. Remove the filter and inspect it (by unscrewing the bluish/purplish cylinder on the back of the chiller). If the filter looks dirty (brownish), replace it with a new one on top of the laser. When we are down to one replacement, order at least 2 new filters from Spectra-Physics/Newport.

    5. Replace the quick disconnect hoses.

    6. Fill the chiller (cap is on top, front right) with milliQ or distilled water borrowed from another lab. DO NOT use deionized water (what we get out of the special faucets in labs), as this will corrode the laser. Ask someone if you are unclear as to what water to use.

    7. Turn on the chiller and make sure it is running. Make sure the set point is 21 degrees C. Make sure the chiller is actually in chilling mode, which means a '-' (minus sign) should be at the far left of the display. If it is not running, press START/STOP to toggle it into running mode. It is critical to the health of the laser that the chiller is actually running.

    8. Wait until the water temperature is within a couple degrees of the target temperature before putting the laser into emission mode.

The chiller filter part is Pentek P5-478, we have ordered from filterfast.com

About every 24 months, one must replace the humidity cartridge / humidity filter / purge filter. BE TRAINED BY SOMEONE WHO KNOWS. To do this:

  1. We need new instructions, it is now in the Power supply

Photos of humidity filter (left) and where it goes on laser (right)

Troubleshooting and notes:

Random notes:

Before contacting tech support over the phone, open the Mai Tai software, open the main panel, the Setup panel, and the Info panel, and take a screen shot.

In the Spectra-Physics software, the pump laser is sometimes referred to as the "Green" laser, and sometimes is referred to as the pump laser.

Wavelengths between 900-1000nm are likely to be harder to obtain (or to work well) if the room is humid. This is because water absorbs light at these wavelengths. If you are having trouble mode-locking at 900-1000nm, try 800nm; this is an easy place to mode-lock at, and then try addressing room or laser humidity.

Indicators to look at that might be useful for Tech support:

P2 and M3 are piezo mirrors; they should be somewhere in the "middle" of their range, not 0, not maximum, but in the middle.

If we get an error "Green Adj Servo Loop error" just wait a few minutes, it may correct itself; (it is just a measure of how close the laser's internal indicators are to those obtained at the factory; the laser usually settles close to these values but an incorrect value may or may not indicate a problem.)

Contacts:

Technical: (use our laser's serial number (now 4571-R1) when you talk to them)

Main number: 1-800-456-2552

If we have trouble getting someone relatively quickly on the 800 number, call down the list of people here:

Our likely support technician:

Jian Liu

jian.liu@spectra-physics.com

Customer Service Engineer

Spectra Physics Lasers

A division of Newport Corp

Cell 408-205-8696

(Mike Ciampa no longer works at SpectraPhysics/Newport)

(Sean Dunegan no longer works at SpectraPhysics/Newport)

Sales:

Paul I. Presunka

PAUL.PRESUNKA@newport.com

Senior Sales Engineer

Spectra-Physics

A Division of Newport Corporation

3635 Peterson Way

Santa Clara, CA 95054

Phone: (508)353-1963

Katheryn Levine (Customer Service) - 408-980-5683

Wilfred Gonzales (Customer Service Engineer, based in New Hampshire) - 603-706-0282

Mike Maikowski (Area Service Coordinator) - 631-484-5689

Warranty contacts:

Tameir Powell (Warranties) - 408-980-4334

Logs:

2019-03-05, Jian Liu replaced laser (old serial 4705, new 4571-R1) so now we have a new laser and the logs below are for the old laser.

On 2013-05-23, Sean Dunegan was out to replace the humidity filter. Humidity was 29% when the cartridge was changed.

2014-07-18, Steve checked humidity filter (fine, less than 4%) and adjusted SHG fine to -73, pump power around 12.00 (11.80-12.00)

2015-06-08, chiller said "Pump Failed". Steve restarted it and it seems to be fine. However, laser humidity is some 23%. Ordered new "purge filter". It is complementary with our service contract. Expect call from Fred with reference 300151623. [Note: this was replaced shortly after and humidity came down.]

2015-August/September

The laser exhibited some difficulty mode-locking. Over the phone, Arani was able to obtain mode-locking by running the laser in "green power mode" (is this really right? was that the way we got it working?)

2015-09-17, Mike Ivy examined the laser, was unable to get it to mode-lock. He ordered a new mode-locking board and replaced it on 2015-09-18 but this did not solve the problem.

2015-09-22 Sean Dunagan returned the laser to its original calibration, made some small tweaks, and the system works okay. It had high humidity (15%) and needs to be recalibrated.

2015-09-23 Steve checked humidity, reads 5% now.

2018-06-18

The laser was extremely humid (21%). Humidity purge filter changed, humidity came down to 4% 2 days later. (This is the right value.)

2018-06-28

The laser exhibited some difficulty mode-locking in the 900 range immediately after warmup. A temporary solution was to tune the laser to 810 or so, let it mode-lock, and then tune up in steps to the 900 range.

On recommendation from Spectra-Physics (service call 300213724), I adjusted the SHG file using the procedure. Here is status before and after the adjustment. The laser now mode-locks to 925nm on initial warm-up, no need to tune to 800 first.

2019-10-31

Mike Ivy updated firmware that should prevent laser from shutting down after each scan.